answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many simple random samples of size 5 can be selected from a population of size 34?

278 256

The number of 5 different item combinations from a pool of 34 different items is

given by:

34C5 = 34!/(5!29!) = 278 256

What is the probability of throwing a 5 on a dice?

1 in 6. U would of thought it was 5/6 but there aren't five 5s on a dice. There's only1.

What does square of probability explain?

If p is the probability that an event will happen once, then the probability that it will happen just twice is p2. The probability it will happen 3 times is p3. The probability it will happen at least once ( ie once or twice or three times ore more times is p + p2 + p3 + ... = p(1-p). For "or" you add probabilities, for "and" you multiply probabilities.

What is the probability of being dealt a spade and heart in a deck of 52 cards?

Number of cards in a deck = 52

Number of cards that are spade = 13

Number of cards that are heart = 13

Probability that the card drawn is a spade and heart =

13/52 + 13/52

= 26/52 or 1/2

When two dice are rolled what is the probability of getting a sum of 5 or 6?

When two dice are rolled, the probability of getting:

2 is 1/36

3 is 2/36

4 is 3/36

5 is 4/36

6 is 5/36

7 is 6/36

8 is 5/36

9 is 4/36

10 is 3/36

11 is 2/36

12 is 1/36

Why is the standard error a smaller numerical value compared to the standard deviation?

Let sigma = standard deviation.

Standard error (of the sample mean) = sigma / square root of (n), where n is the sample size. Since you are dividing the standard deviation by a positive number greater than 1, the standard error is always smaller than the standard deviation.

What does it mean when the standard error value is smaller than the standard deviation?

It simply means that you have a sample with a smaller variation than the population itself. In the case of random sample, it is possible.

What is the probability that you will get a 3-digit combination correct if no digits in the combination can be repeated?

If the digit 0 is allowed then the answer is 1/(10*9*8) = 1/720 = 0.0014 approx.

If 0 is disallowed, then 1/(9*8*7) = 1/504 = 0.0020 approx.

What is the probability of flipping a coin three times and it landing on heads all three times?

1/8.

The probability of flipping a coin three times and it landing on head is 1/2, as a coin only has two sides.

You flip a coin three times, therefore the answer is (1/2)^3 = 1/8.

When do you use poisson probability?

When you have independent events which have a constant probability of occurrence over an interval of space or time.

What is the probability of the next four births at a hospital being all girls?

Assuming that the births at the hospital are equally likely to be of either gender then the answer is (1/2)4 = 1/16

Can probability of an event be negative?

An event will happen or not so its probabilty is between 0 (never happen) and 1 (will surely happen).

The probability is a measure of its chance to happen.

A negative probability would have no sense.

You can think about probabilty this way: imagine a 1-litre bottle, it can be empty (0 litre) or full (1 litre) and any value between 0 and 1 but the value can't be negative.

How many ways can the word Tennessee be arranged?

Well, there are 9 letters in Tennessee, and, based on the question, you can only use each letter once.

The first letter has 9 possible outcomes.

The second letter only has 8, because you have already used one.

The third would have 7, etc.

9 * 8 * 7 * 6 * 5 * 4 * 3 * 2 * 1 = 322,880

Trending Questions
How do you measure probability? A sack contains red and blue marbles The ratio of red marbles to blue marbles is 4 to 3 If there are 16 red marbles in the sack how many blue marbles are in the sack? If the total area of a dartboard is 40000 mm2 and the area of the second ring is 17500 mm2 what is the probability of landing in that second ring? What are the chances of getting hit by a tree? What are the chances of a family having five boys in a row? What are the odds of 2 dice totaling a 7 on one throw? How many outcomes are there for the event of rolling a 6 sided number cube and then spinning a 10 part spinner? Frequency approach for probability? When did the study of ergonomics begin? What is the total number of outcomes when rolling two number cubes and tossing one coin? Does A probability near 0 indicates an unlikely event.? What is the probability of throwing a six on a dice? What probability approach can be applied only to experiment with equally likely outcomes? Probability of picking one red card of a deck of 52 playing cards? There are 10 marbles in a bag 5 are blue 2 are yellow and 3 are purple what is the probability of picking a purple marble? Is the gender of college students a discrete random variable a continuous random variable or not a random variable? What is Keynes's theory on probability? What is the probability a prime number will show up on a number cube? How do you find the area to the right of the z-score? Why was the X and Y letters used to refer to male and female chromosomes?